# Bayes2IMC: In-Memory Computing for Bayesian Binary Neural Networks

### v1

@@ V Conclusion
Through a hardware-software co-optimization flow, output logits were corrected to enhance accuracy and the expected calibration error. Additionally, the reparametrization technique supports a straightforward drift compensation method that fully mitigates drift-induced accuracy degradation. We also successfully demonstrated uncertainty quantification and decomposition into aleatoric and epistemic uncertainties using Bayes2IMC.
Our hardware projections, informed by literature and CAD projections, indicate that while Bayes2IMC operates 28×2-8$\times$2 - 8 × slower than an SRAM-based architecture, it achieves up to 10×10$\times$10 × higher efficiency in terms of GOPS/W/mm2. We also projected that Bayes2IMC is up to $20\%$ more power-efficient compared to the state-of-the-art. Our analysis reveals that the peripheral circuitry, both digital and analog, contributes significantly more to area and power consumption than the crossbar or memory elements themselves. Future efforts to improve efficiency must thus focus on optimizing peripheral circuits.
The techniques, algorithms, and co-optimization framework discussed in this paper are demonstrated with PCM devices but are versatile enough to be applied to other noisy NVM device-based compute architectures for implementing binary Bayesian networks.

### v2

@@ V Conclusion
Through a hardware-software co-optimization flow, output logits were corrected to enhance accuracy and the ECE. Additionally, the reparametrization technique supports a straightforward drift compensation method that fully mitigates drift-induced accuracy degradation. We also successfully demonstrated uncertainty quantification and decomposition into aleatoric and epistemic uncertainties using Bayes2IMC.
Our hardware projections, informed by literature and CAD projections, indicate that while Bayes2IMC operates 28×2-8$\times$2 - 8 × slower than an SRAM-based architecture, it achieves up to 10×10$\times$10 × higher efficiency in terms of GOPS/W/mm2. We also projected that Bayes2IMC is up to $20\%$ more power-efficient than the state-of-the-art. Our analysis reveals that the digital and analog peripheral circuitery contributes significantly more to area and power consumption than the crossbar or memory elements themselves. Future efforts to improve efficiency must thus focus on optimizing peripheral circuits. The techniques, algorithms, and co-optimization framework discussed in this paper are demonstrated with PCM devices but are versatile enough to be applied to other noisy NVM device-based compute architectures for implementing binary Bayesian networks.